Provide the missing data in the following table for a distributor of martial arts products. (Enter Turnover and ROI answe
Answer #1

Margin=Net operating income/Sales

Turnover=Sales/Average operating assets

ROI=Net operating income/Average operating assets

Alpha Bravo Charlie
Sales (457,000*5)=$2,285,000 307500 (50040/0.09)=$556,000
Net operating income (2,285,000*5%)=$114250 30750 50040
Average operating assets 457,000 (30750/0.25)=$123,000 (50040/0.18)=$278,000
Margin 5% (30750/307500)=10% 9%
Turnover 5 (307500/123000)=2.5 (556,000/278,000)=2
ROI (114250/457000)=25% 25% 18%